OpenStudy (anonymous):

Suppose Y~p(y)=1/30(y-5)^2 for y = 1,2,3,4. Calculate E(1/Y).

OpenStudy (anonymous):

\[\frac{ 1 }{30 }(y-5)^2 = p(y)\]